# TathyaNyaya and FactLegalLlama: Advancing Factual Judgment Prediction and Explanation in the Indian Legal Context

### v1

@@ Abstract
In the landscape of Fact-based Judgment Prediction and Explanation (FJPE), reliance on factual data is essential for developing robust and realistic AI-driven decision-making tools. This paper introduces TathyaNyaya, the largest annotated dataset for FJPE tailored to the Indian legal context, encompassing judgments from the Supreme Court of India and various High Courts. Derived from the Hindi terms “Tathya" (fact) and “Nyaya" (justice), the TathyaNyaya dataset is uniquely designed to focus on factual statements rather than complete legal texts, reflecting real-world judicial processes where factual data drives outcomes. Complementing this dataset, we present FactLegalLlama, an instruction-tuned variant of the LLaMa-3-8B Large Language Model (LLM), optimized for generating high-quality explanations in FJPE tasks. Fine-tuned on the factual data in TathyaNyaya, FactLegalLlama integrates predictive accuracy with coherent, contextually relevant explanations, addressing the critical need for transparency and interpretability in AI-assisted legal systems. Our methodology combines transformers for binary judgment prediction with FactLegalLlama for explanation generation, creating a robust framework for advancing FJPE in the Indian legal domain. TathyaNyaya not only surpasses existing datasets in scale and diversity but also establishes a benchmark for building explainable AI systems in legal analysis. The findings underscore the importance of factual precision and domain-specific tuning in enhancing predictive performance and interpretability, positioning TathyaNyaya and FactLegalLlama as foundational resources for AI-assisted legal decision-making.
@@
Shubham Kumar Nigam1   Balaramamahanthi Deepak Patnaik1   Shivam Mishra1 Noel Shallum3   Kripabandhu Ghosh2   Arnab Bhattacharya1 1 IIT Kanpur, India  2 IISER Kolkata, India  3 Symbiosis Law School Pune, India {sknigam, deepak, shivammishra, arnabb}@cse.iitk.ac.in kripaghosh@iiserkol.ac.in,noelshallum@gmail.com
@@ 1 Introduction
The integration of AI technologies into the legal domain holds immense potential for improving the efficiency, accessibility, and transparency of judicial processes, especially in jurisdictions like India, where case backlogs place a significant burden on the courts. Among the emerging solutions, Fact-based Judgment Prediction and Explanation (FJPE) aims to predict judicial outcomes and provide explainable rationales by focusing exclusively on the factual elements of a case. By highlighting the core facts rather than relying on entire legal documents, FJPE seeks to deliver more realistic early-phase predictions, akin to conditions when a judge must form an opinion before hearing all arguments or awaiting additional documents.
Recent studies have explored FJPE by extracting factual elements from full case texts or summarizing multiple components like statutes, ratios of decisions, and arguments Nigam et al. (2024b); Nigam and Deroy (2024). While these efforts represent a step toward fact-centric approaches, they often rely on limited or automatically extracted data that may not always be reliable. This distinction is crucial: unlike Court Judgment Prediction with Explanation (CJPE), which may utilize entire judgments including judicial reasoning and statutory references, FJPE restricts itself to facts presented at the time of decision-making. Such a focus mimics the constraints of real-world legal reasoning, where judges form preliminary opinions based on presented facts before considering lengthy arguments, precedents, or statutes.
In this paper, we introduce TathyaNyaya, the first large-scale, expertly annotated dataset dedicated to FJPE in the Indian legal context. Derived from the Hindi terms "Tathya" (fact) and "Nyaya" (justice), TathyaNyaya places factual segments at the heart of predictive modeling, thereby providing a robust foundation that eschews the noise and complexity of full legal documents. This annotated dataset addresses a critical gap: while prior work has attempted fact-based prediction using incomplete or non-annotated sources, TathyaNyaya represents the first systematic effort to assemble, annotate, and align factual information with judicial outcomes and explanations. By simplifying the input space to factual data, our dataset enables models to more accurately reflect real-world conditions, offer transparent and early-stage predictive insights, and reduce reliance on domain-level heuristics or post-hoc explanations.
TathyaNyaya incorporates judgments from the Supreme Court of India (SCI) and various High Courts, ensuring broad coverage and diversity. Comprising four components—NyayaFacts, NyayaScrape, NyayaSimplify, and NyayaFilter, the dataset caters to different facets of factual extraction, simplification, and retrieval, thereby facilitating comprehensive FJPE research.
To complement this dataset, we present FactLegalLlama, an instruction-tuned variant of LLaMa-3-8B fine-tuned specifically for FJPE tasks. FactLegalLlama is explicitly adapted to produce faithful, fact-grounded explanations.
Our key contributions are summarized as:
@@
  • TathyaNyaya Dataset: We introduce TathyaNyaya, the first extensively annotated, purely fact-centric dataset for judgment prediction and explanation in the Indian legal context, providing a stable benchmark for FJPE research. The dataset is structured into four components, each addressing distinct aspects of factual legal analysis.
  • Factual Basis for Prediction: Unlike previous datasets that use full legal texts, TathyaNyaya focuses on factual information, allowing for more realistic and targeted legal judgment predictions.
  • FactLegalLlama for Explanation: We introduce FactLegalLlama, a LLaMa-3-8B model fine-tuned to generate high-quality, factual explanations for judicial outcomes, enhancing interpretability and trust in AI-assisted legal systems.
To ensure reproducibility and encourage further research, the dataset and model code will be made publicly available soon.
@@ 2 Related Work
The domain of Legal Judgment Prediction (LJP) has seen substantial progress, driven by a need for efficient and transparent decision-making aids in the legal system. Early foundational works Aletras et al. (2016); Chalkidis et al. (2019); Feng et al. (2021) focused on predicting case outcomes while emphasizing interpretability, laying the groundwork for methodologies and benchmark datasets like CAIL2018 Xiao et al. (2018) and ECHR-CASES Chalkidis et al. (2019). These resources fostered the development of advanced models including TopJudge and MLCP-NLN, yet a consistent performance gap remains between automated predictions and actual judicial decisions.
Within the Indian legal ecosystem, researchers have introduced datasets such as ILDC Malik et al. (2021) and PredEx Nigam et al. (2024a), alongside related efforts Nigam et al. (2022); Malik et al. (2022); Nigam et al. (2023), underscoring the significance of domain-specific datasets and the need for explanations that can be integrated into real-world legal workflows. Studies leveraging Large Language Models (LLMs) Vats et al. (2023); Nigam et al. (2024a) have shown that models such as GPT-3.5 Turbo and LLaMa-2 can adapt to Indian legal texts, further diversifying approaches to LJP and interoperability. Fact-based judgment prediction has gained prominence as a realistic approach to LJP, focusing on predictions derived from case facts rather than full case judgments. Nigam et al. (2024b) and Nigam and Deroy (2024) explore LJP based on facts, arguing that this approach better simulates real-world scenarios.
Cross-jurisdictional research, including that by Zhao et al. (2018), has extended LJP methodologies across different legal frameworks. Multilingual efforts like Niklaus et al. (2021); Kapoor et al. (2022) and approaches integrating event extraction and multi-stage reasoning Feng et al. (2022) continue to broaden the capabilities of LJP systems.
@@ 3 Task Description
The Fact-based Judgment Prediction and Explanation (FJPE) task consists of two subtasks:
Task A: Judgment Prediction: This is a binary classification problem. Given the factual information of the legal judgment as input, the objective is to predict whether the decision favors or goes against the appellant. The prediction is represented by binary labels: "1" indicates that the appeal is accepted (i.e., if any part of the appeal is accepted, the decision is considered in favor of the appellant), while "0" indicates that the appeal is rejected.
Task B: Rationale Explanation: This subtask involves generating an explanation or rationale that justifies the predicted decision based on the provided factual information of the case. The goal is to provide a clear understanding of the reasoning behind the predicted outcome, grounded in the facts presented.
Figure 2 in the Appendix illustrates the overall process of fact-based judgment prediction and explanation employed in our study, outlining the sequential steps from prediction to explanation based on the factual data provided.
@@ 4 Dataset
In this research, we introduce TathyaNyaya, a comprehensive dataset explicitly designed for Fact-based Judgment Prediction and Explanation (FJPE) in the Indian legal domain. This dataset consists of four distinct components: (1) NyayaFacts—expert-annotated data that serves as the gold standard for prediction and explanation tasks, (2) NyayaScrape—automated fact-extracted data obtained through machine-driven processes, (3) NyayaSimplify—a user-friendly dataset created by paraphrasing complex legal language, and (4) NyayaFilter—a binary fact vs. non-fact classification dataset designed to streamline the retrieval of relevant factual information. Together, these components form the largest and most diverse factual dataset in the Indian judiciary, enabling the development and evaluation of advanced AI models for transparent and interpretable judgment prediction and explanation. By focusing exclusively on factual data, TathyaNyaya addresses a critical gap in the field, paving the way for more robust and realistic AI-driven solutions tailored to the Indian legal context.
Figure 1 illustrates the TathyaNyaya dataset creation pipeline. It provides a high-level overview of how each component which is derived, from expert-curated facts and machine-driven extraction, to fact segmentation and paraphrasing. This end-to-end pipeline ensures that the final dataset captures both breadth and depth in factual legal information, supporting the FJPE task.
@@
4.1 Dataset Compilation and Statistics
4.1.1 NyayaFacts
NyayaFacts comprises a subset of Supreme Court of India (SCI) and High Court judgments carefully annotated by legal experts. These annotations highlight key factual segments that significantly influence judicial outcomes, serving as high-quality ground truth for both judgment prediction and rationale explanation. After refining and preprocessing, this subset serves as the gold standard for evaluating prediction and explanation tasks.
In particular, the validation and test data were derived from the NyayaFacts Single subset to maintain consistency during evaluation, while the training data include both single and multi-case judgments, offering a broad learning landscape. Table 1 provides comprehensive statistics. NyayaFacts thus provides a high-quality benchmark for both judgment prediction and explanation tasks.
@@ 4.1.3 NyayaSimplify
NyayaSimplify focuses on making complex legal texts more accessible by paraphrasing the NyayaFacts test data into clearer, more concise language. Using LLaMA-3-70B-Instruct, intricate legal jargon is transformed into user-friendly text without altering the factual content or legal reasoning. Since NyayaSimplify is derived directly from NyayaFacts, the majority of dataset statistics remain comparable, with the primary difference being a reduced average word count. This simplification aims to improve both the accuracy and interpretability of models in FJPE tasks. Detailed prompt used for paraphrasing is provided in Appendix Table 7.
@@ 4.1.4 NyayaFilter
# Documents 14,134 1,197 2,389
# Sentences 15,83,858 49,671 56,240
Avg # Words 29.03 29.00 34.00
Avg # Facts/Document (%) 25.1 20.61 22.7
Overall Facts (%) 20.34 12.86 18.46
@@
# Documents 14,134 1,197 2,389
# Sentences 4,04,349 3,36,478 248,433
Avg # Words 28.06 27.00 30.00
Avg # Non-Facts/Document (%) 74.9 79.39 77.3
Overall Non-Facts (%) 79.66 87.14 81.54
@@
4.2 Annotation Methodology and Quality Assurance
@@ 4.2.4 Quality Control Framework
This robust framework ensured the annotations’ accuracy and credibility, making NyayaFacts a reliable benchmark for evaluating prediction and explanation models.
@@ 5 Methodology
In this section, we present our overall methodology for extracting factual segments from legal judgments, training our custom model FactLegalLlama for Fact-based Judgment Prediction and Explanation (FJPE), and finally addressing both the prediction-only and prediction-with-explanation tasks. We also detail the prompts we used and instruction-tuning strategies employed to refine our model’s outputs.
5.1 Fact Extraction from Full Legal Judgments
To prepare the dataset for Fact-based Judgment Prediction and Explanation (FJPE), we first extracted the factual statements from full-text legal judgments. We adopted a streamlined binary classification approach by fine-tuning a BiLSTM-CRF model Ghosh and Wyner (2019), a previous state-of-the-art (SoTA) model for semantic segmentation of legal documents. Instead of using the original multi-class rhetorical role framework, which distinguishes between roles such as issue, statute, precedent, and argument, we simplified the task by treating all non-factual segments as a single class labeled "non-facts."
This transformation into a binary classification problem enabled the model to focus solely on identifying factual segments critical to judgment prediction. Training was conducted using the NyayaFacts multi, which provided expert-annotated labels for factual and non-factual segments. By isolating the facts, we laid the groundwork for developing AI models capable of making decisions and generating explanations based solely on factual data. This preprocessing ensured that the subsequent models trained on the dataset remained focused on the most relevant and actionable information in legal cases.

### v2

@@ Abstract
In the legal domain, Fact-based Judgment Prediction and Explanation (FJPE) aims to predict judicial outcomes and generate grounded explanations using only factual information, mirroring early-phase legal reasoning. Motivated by the overwhelming case backlog in the Indian judiciary, we introduce TathyaNyaya, the first large-scale, expert-annotated dataset for FJPE in the Indian context. Covering judgments from the Supreme Court and multiple High Courts, the dataset comprises four complementary components, NyayaFacts, NyayaScrape, NyayaSimplify, and NyayaFilter, that facilitate diverse factual modeling strategies. Alongside, we present FactLegalLlama, an instruction-tuned LLaMa-3-8B model fine-tuned to generate faithful, fact-grounded explanations. While FactLegalLlama trails transformer baselines in raw prediction accuracy, it excels in generating interpretable explanations, as validated by both automatic metrics and legal expert evaluation. Our findings show that fact-only inputs and preprocessing techniques like text simplification and fact filtering can improve both interpretability and predictive performance. Together, TathyaNyaya and FactLegalLlama establish a robust foundation for realistic, transparent, and trustworthy AI applications in the Indian legal system.
@@
Shubham Kumar Nigam1,4∗†   Balaramamahanthi Deepak Patnaik1∗ Shivam Mishra1Noel Shallum2Kripabandhu Ghosh3Arnab Bhattacharya1 1IIT Kanpur, India  3Symbiosis Law School Pune, India  2IISER Kolkata, India 4University of Birmingham, Dubai, United Arab Emirates {shubhamkumarnigam, bdeepakpatnaik2002, shivam1602m, noelshallum} @gmail.com kripaghosh@iiserkol.ac.in,arnabb@cse.iitk.ac.in
[…]
@@ 1 Introduction
The integration of AI technologies into the legal domain holds immense potential for improving the efficiency, accessibility, and transparency of judicial processes, particularly in countries like India, where case backlogs severely burden the courts. As of recent estimates, over 50 million cases are pending across various courts in India (njdc-district), resulting in delays that can stretch into decades. In this context, early-phase legal decision support, i.e., prediction based solely on factual information available at the beginning of a case, has emerged as a highly relevant research goal. Among the emerging solutions, Fact-based Judgment Prediction and Explanation (FJPE) offers a promising direction. FJPE aims to predict judicial outcomes and provide rationales using only the factual elements of a case, without relying on arguments, precedents, or judicial reasoning. This mirrors real-world scenarios where stakeholders, judges, lawyers, or litigants, must assess case strength based on initial facts to decide whether to proceed, allocate resources, or pursue alternative legal remedies. Furthermore, factual records are often the most reliably documented and readily available components in early legal proceedings, especially in resource-constrained environments.
While previous studies have attempted fact-centric modeling by summarizing multiple legal components or relying on automatically extracted facts nigam-etal-2024-rethinking; 10.1145/3632754.3632765, these approaches often lack reliable ground truth and blur the boundaries between pure factual inputs and broader legal discourse. Moreover, such works typically reference the full case context, statutes, or reasoning, placing them closer to the domain of CJPE, which includes post-filing evidence and legal argumentation. In contrast, FJPE distinctly isolates factual segments to simulate the setting of early-phase legal reasoning, where preliminary decisions may be formed even before formal hearings begin. To advance this direction, we introduce TathyaNyaya, the first large-scale, expertly annotated dataset explicitly designed for FJPE in the Indian legal context. The term combines the Hindi words “Tathya” (fact) and “Nyaya” (justice), underscoring its foundation in factual legal analysis. Unlike prior datasets, TathyaNyaya does not rely on heuristics or summarization techniques; instead, it offers cleanly annotated factual inputs aligned with judicial outcomes and explanations, allowing for reproducible, interpretable, and practical early-stage prediction models.
TathyaNyaya comprises judgments from the Supreme Court of India (SCI) and various High Courts and is organized into four components: NyayaFacts, NyayaScrape, NyayaSimplify, and NyayaFilter. These components support a wide range of fact-centric tasks, from expert annotations and simplified factual paraphrasing to fact vs. non-fact segmentation. Complementing the dataset, we introduce FactLegalLlama, an instruction-tuned version of LLaMa-3-8B, fine-tuned on TathyaNyaya to perform FJPE tasks. While transformer-based models are strong in predictive performance, FactLegalLlama demonstrates the ability to generate faithful and interpretable factual explanations, thus bridging predictive modeling with legal reasoning.
Our key contributions are:
@@
  • TathyaNyaya Dataset: We introduce the first extensively annotated, purely fact-centric dataset for judgment prediction and explanation in the Indian legal domain, structured into four components tailored for factual segmentation, simplification, and retrieval.
  • Early-Phase Legal Reasoning: We focus on realistic and societally impactful early-phase decision-making settings where predictions are made using only the facts, reflecting constraints and needs of India’s overburdened judiciary.
  • FactLegalLlama for Explanation: We propose FactLegalLlama, an instruction-tuned LLaMa-3-8B model designed to generate faithful and fact-grounded explanations for judicial outcomes. It excels in producing coherent and semantically aligned rationales.
To foster transparency and reproducibility, we have implemented our data set, code, and model through a GitHub repository [^1] TathyaNyaya-and-FactLegalLlama GitHub Repo.
@@ 2 Related Work
Judgment Prediction has evolved into one of the most studied problems in legal NLP, driven by the need for computational assistance in judicial decision-making. Foundational studies such as aletras2016predicting, chalkidis2019neural, and feng2021recommending first demonstrated the feasibility of predicting legal outcomes from textual case records, leading to benchmark datasets such as CAIL2018 xiao2018cail2018 and ECHR-CASES chalkidis2019neural. These works inspired architectures that integrate predictive performance with interpretability, setting the foundation for subsequent developments.
In the Indian context, significant efforts have been made to address judgment prediction and related downstream tasks. The ILDC corpus malik-etal-2021-ildc provided the first large-scale annotated dataset for CJPE, while later initiatives such as NyayaAnumana nigam2024nyayaanumana and PredEx nigam2024legaljudgmentreimaginedpredex expanded interpretability by combining factual and reasoning-based perspectives. Subsequent research explored complementary directions, including Legal Question Answering (AILQA) nigam2023legal, rhetorical role segmentation for legal document structuring malik2021semantic; nigam2025legalseg, and structured document generation with VidhikDastaavej nigam2025structured. Further, vats-etal-2023-llms; nigam2022nigam analyzed the efficacy of Large Language Models (LLMs) such as GPT and LLaMA for statute and judgment prediction, exposing both potential and bias in legal AI. Recent datasets like NyayaRAG nigam2025nyayarag, and IBPS srivastava2025ibps continue advancing realistic, explainable, and fact-grounded AI frameworks within the Indian judicial system.
Fact-based LJP has emerged as a more interpretable and realistic paradigm, focusing on early-phase factual understanding. Works such as nigam-etal-2024-rethinking and 10.1145/3632754.3632765 explicitly emphasized the challenge of reasoning solely from facts while maintaining legal coherence. However, those approaches relied on extractive or summarization-based proxies for factual representation. Our work directly addresses this gap by providing expert-annotated factual data in NyayaFacts and by fine-tuning FactLegalLlama to generate faithful, fact-grounded explanations. This focus aligns with the broader explainability and interpretability movement in legal NLP, where models are expected to provide transparent, human-understandable justifications marino2023automatic; santosh2024hiculr.
Beyond India, multilingual and cross-jurisdictional research has broadened the global reach of LJP. The SwissJudgmentPrediction dataset niklaus2021swiss and HLDC corpus kapoor-etal-2022-hldc extended analysis to non-English and Hindi legal texts, respectively. Similarly, rhetorical and event-based modeling approaches feng-etal-2022-legal; santosh2025lecopcrlegalconceptguidedprior enriched semantic structure understanding. Parallel initiatives like the SAIL Symposium ghosh2023report; ganguly2023legal have consolidated research in Legal IR, information retrieval, and reasoning tasks.
@@ 3 Task Description
We divide TathyaNyaya documents into 2 sets:
  • Single: Either it contains a single petition or multiple petitions where all decisions are identical.
  • Multi: It contains multiple appeals with different outcomes. For simplicity, we convert all partially accepted cases into accepted, preserving the binary classification setup. Thus, both single and multi datasets support binary classification.
The task consists of two subtasks:
Task A: Judgment Prediction: This is a binary classification problem. Given the factual information of a legal case, the goal is to predict whether the judgment favors the appellant or not. A label of "1" denotes acceptance (including partially accepted cases), and "0" denotes complete rejection.
Task B: Rationale Explanation: This subtask involves generating a textual explanation for the predicted decision. The rationale should be grounded in the provided factual information and reflect the reasoning that supports the outcome.
Figure 2 in the Appendix illustrates the overall FJPE pipeline, outlining the stages from fact input to prediction and explanation generation.
@@ 4 Dataset
In this research, we introduce TathyaNyaya, a comprehensive dataset explicitly designed for FJPE in the Indian legal domain. This dataset consists of four distinct components: (1) NyayaFacts: expert-annotated data that serves as the gold standard for prediction and explanation tasks, (2) NyayaScrape: automated fact-extracted data obtained through machine-driven processes, (3) NyayaSimplify: a user-friendly dataset created by paraphrasing complex legal language, and (4) NyayaFilter: a binary fact vs. non-fact classification dataset designed to streamline the retrieval of relevant factual information. Together, these components form the largest and most diverse factual dataset in the Indian judiciary, enabling the development and evaluation of advanced AI models for transparent and interpretable judgment prediction and explanation. By focusing exclusively on factual data, TathyaNyaya addresses a critical gap in the field, paving the way for more robust and realistic AI-driven solutions tailored to the Indian legal context.
Figure 1 illustrates the TathyaNyaya dataset creation pipeline. It provides a high-level overview of how each component which is derived, from expert-curated facts and machine-driven extraction, to fact segmentation and paraphrasing. This end-to-end pipeline ensures that the final dataset captures both breadth and depth in factual legal information.
4.1 Dataset Compilation and Statistics
4.1.1 NyayaFacts
NyayaFacts comprises a subset of different Court judgments carefully annotated by legal experts. These annotations highlight key factual segments that significantly influence judicial outcomes, serving as high-quality ground truth for both judgment prediction and rationale explanation. After refining and preprocessing, this subset serves as the gold standard for the FJPE task.
In particular, the validation and test data were derived from the NyayaFacts Single subset to maintain consistency during evaluation, while the training data include both single and multi-case judgments, offering a broad learning landscape. Table 1 provides comprehensive statistics.
@@ 4.1.3 NyayaSimplify
NyayaSimplify aims to enhance model performance and interpretability by transforming complex legal texts into simplified, paraphrased versions. Since most LLMs are pre-trained on general-purpose corpora and not on legal-specific jargon, they often struggle with the dense and domain-specific language found in court judgments. To address this, we paraphrased the NyayaFacts test data using the LLaMA-3-70B-Instruct model. This transformation preserves the factual and legal integrity of the original content while expressing it in more accessible, human-readable language.
The resulting dataset allows us to evaluate whether simplifying legal language helps general-purpose models better understand and reason about legal facts. While most dataset statistics remain consistent with NyayaFacts, the average word count is notably reduced, indicating a successful simplification. Our findings suggest that simplification improves both the accuracy and interpretability of models on FJPE tasks. Prompt template used for paraphrasing is included in Appendix Table 9.
@@ 4.1.4 NyayaFilter
# Documents 13,629 1,197 2,389
# Sentences 3,62,658 30,561 56,240
Avg # Words 29.00 29.00 34.00
Avg # Facts/Document (%) 23.6 23.03 22.7
Overall Facts (%) 19.16 19.09 18.46
@@
# Documents 13,629 1,197 2,389
# Sentences 15,29,998 1,29,543 2,48,433
Avg # Words 28.00 28.00 30.00
Avg # Non-Facts/Document (%) 76.4 76.97 77.3
Overall Non-Facts (%) 80.84 80.91 81.54
@@
4.2 Annotation Process & Quality Assurance
@@ 4.2.4 Quality Control Framework
  • Framework Evaluation and Reproducibility Assessment: To assess the stability and reproducibility of the factual annotation process, we conducted an independent evaluation involving two additional legal annotators. A randomly selected subset of 50 documents from the NyayaFacts dataset was re-annotated using the same annotation guidelines. The original annotations were treated as the reference (gold), and textual similarity metrics, BLEU, METEOR, ROUGE-1/2/L, and BERTScore were computed to compare the new annotations with the original.
    Annotator Pair BLEU METEOR R1 R2 RL BERTScore
    A2 vs A1 (Gold) 0.78 0.62 0.85 0.68 0.80 0.91
    A3 vs A1 (Gold) 0.75 0.60 0.83 0.65 0.78 0.89
    Table 3: Agreement scores between original annotations (A1) and two additional annotators (A2 and A3) across 50 re-annotated legal cases.
    These results demonstrate strong agreement between the independent annotators and the original gold annotations. In particular, high ROUGE-1 (>>0.80) and BERTScore (>>0.89) values indicate consistent extraction of factually relevant segments, both lexically and semantically. This empirical evidence supports the reproducibility and reliability of the annotation framework employed for NyayaFacts, validating its use as a high-quality benchmark for fact-based legal judgment prediction and explanation.
@@ 5 Methodology
In this section, we present our overall methodology for extracting factual segments from legal judgments, training our custom model FactLegalLlama for FJPE, and finally addressing both the prediction-only and prediction-with-explanation tasks. We also detail the prompts we used and instruction-tuning strategies employed to refine our model’s outputs.
5.1 Fact Extraction from Full Judgments
To prepare the dataset for FJPE, we first extracted the factual statements from full-text legal judgments. We adopted a streamlined binary classification approach by fine-tuning a BiLSTM-CRF model ghosh2019identification, a previous state-of-the-art (SoTA) model for semantic segmentation of legal documents. Instead of using the original multi-class rhetorical role framework, which distinguishes between roles such as issue, statute, precedent, and argument, we simplified the task by treating all non-factual segments as a single class labeled "non-facts". This transformation into a binary classification problem enabled the model to focus solely on identifying factual segments critical to judgment prediction. Training was conducted using the NyayaFacts multi, which provided expert-annotated labels for factual and non-factual segments. By isolating the facts, we laid the groundwork for developing AI models capable of making decisions and generating explanations based solely on factual data. This preprocessing ensured that the subsequent models trained on the dataset remained focused on the most relevant and actionable information in legal cases.